OpenStudy (anonymous):

Julianne wants to take a trip around the world. she plans to deposit $175 at the beginning of ach month into an investment with a 3.25% interest rate, compounded monthly. how much will she have in the account after 12 years?

OpenStudy (jdoe0001):

\(\Large { A=P\left(1+\frac{r}{n}\right)^{nt} } \\ \quad \\ A=\textit{current amount}\\ P=\textit{original amount deposited}\to \$175\\ r=rate\to 3.25\%\to \frac{3.25}{100}\to 0.0325\\ n=\textit{times it compounds per year, 12 months per year}\to 12\\ t=years\to 12 \)